A COMPUTATIONAL STUDY OF DOUBLE PROTON TRANSFER IN THE TROPOLONE-HF COMPLEX

\underline{LORI A. BURNS}, DANIEL MURDOCK, and PATRICK H. VACCARO, Department of Chemistry, Yale University, P.O. Box 208107, New Haven, CT 06520-8107.

Many processes of chemical and biological importance depend critically upon the cooperative action of several hydrogen bonds, as do related multiple proton-transfer (MPT) events. The distinct hydrons participating in MPT pathways interact not only with restricted skeletal motions of the molecular framework, but also with the large-amplitude displacements of other shuttling atoms. Consequently, the dynamics exhibited by these processes are expected to be more complex and varied than those of their single-proton counterparts, such as the prototypical tropolone (TrOH) system. This talk will explore the binary cleft-complex (TrOH-HF) formed by binding HF into the tropolone reaction site, a model adduct for probing the sequential versus concerted nature of double proton-transfer processes. The ground electronic state of TrOH-HF shares many distinguishing features inherent to bare TrOH, including a symmetric double-minimum potential well in which two stable and equivalent structures of planar C_{s} symmetry are separated by a C_{2v} transition state. Our ab initio studies have utilized the coupled cluster ansatz (CCSD/EOM–CCSD) in conjunction with extensive correlation-consistent basis sets to examine the X̃^{1}A_{1} and Ã^{1}B_{2} electronic manifolds. The synchronous double protontransfer reaction in TrOH-HF is predicted to exhibit a barrier of 2307.2 cm^{−1} height, representing a 30% drop from the analogous value for TrOH, 3294.2 cm^{−1}. Redistribution of charge density upon π^{∗} ← π electronic excitation transforms the potential energy surface, yielding diminished barriers for both the bare (1135.8 cm^{−1}) and complexed (945.8 cm^{−1}) species. Intriguingly, the pathway for double proton transfer in electronically-excited TrOH-HF shows evidence of nonplanarity, including the the presence of a twisted (C_{2}) transition-state geometry in which the lone imaginary frequency combines concerted O–H/F–H stretching motions with a geared rotation of the reaction site. The potential energy landscapes of TrOH and TrOH-HF, as well as the underlying nature of the vibrationally-mediated hydron migration process, will be discussed within the framework of the encompassing G_{4} molecular symmetry group.